OpenStudy (anonymous):

The volume of a box is given the formula v=lwh. The of the box is 320. Find the width and length.

OpenStudy (anonymous):

You have bunch of solution. :)

OpenStudy (anonymous):

For example: V =320 and 320 = 1*10*32, I can say l =1, w =10, h =32, right? or 320 = 2*5*32, then l =2, w = 5 and h =32 so on and so on.....